3. Fill in the blanks in the table below. Assume that MPC is constant for everyone in the economy.

MPC

Spending Multiplier

Change in government spending

Change in income

100

$15

20

$2,000

0.6

-$400

0.5

$450

$900
Answer #1
MPC Spending multiplier=1/1-MPC Change in government spending Change in income
1-1/100 = 0.99 100 15 (100)(15)=1500
1-1/20 = 0.95 20 (2000/20) = 100 2000
0.6 1/1-0.6 = 2.5 -400 (-400)(2.5) = 1000
0.5 1/1-0.5 = 2 450 (450)(2) =900